# The Impact of Copyrighted Material on Large Language Models: A Norwegian Perspective

### v1

@@ 6 Results
The evaluation of the trained models demonstrated that incorporating copyrighted materials provided a measurable boost in performance across a range of NLP tasks. To illustrate the overall performance differences, Figure 2 shows the total scores across all skills, averaged by task for each model. Non-aggregated scores for all tasks and models are available at the Mímir repository. [^13] https://github.com/mimir-project/mimir-evaluation-suite
@@ 6.1 Core Models
As shown in Table 4 and Figure 3, the performance analysis of core models across various tasks reveals distinct strengths for different configurations. The base (warm-started) configuration consistently excels in Sentiment Analysis, World Knowledge, and Norwegian Language. In contrast, the extended (warm-started) configuration leads in Fairness & Truthfulness, Reading Comprehension, Commonsense Reasoning, Translation, and Variation & Readability, indicating its robust performance for language-intensive tasks. The base configuration generally lags behind others, scoring the lowest across multiple tasks. Meanwhile, the extended configuration performs well, particularly in Summarization. Furthermore, it indicates that we could leverage the existing metadata available at the National Library to tailor subsets of the copyrighted material and build models that excel at specific tasks. However, the difference between the base and extended warm-started models is very small. Further testing is required to assess whether this difference is still statistically significant.
baseextendedbase(warm)extended(warm)Mistral 7Bv0.1448.01482.98480.28441.83413.98Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability
Figure 3: Cumulative score of the core pre-trained models under different regimes. Including original Mistral 7B v0.1 for reference.
@@ 6.2 Domain-tuned Models
To further explore the specific effects of different types of training data, we analyzed the gains in performance by focusing on different sub-corpora, such as newspapers, books, and mixed datasets. Figure 4 provides an overview of the average percentage gains for models trained on various data configurations compared to the base model. It shows that the extended model exhibits the highest average gain at 6.73%, indicating substantial overall improvement. The addition of nonfiction books and newspapers follows with a 6.52% gain, and the addition of only newspapers shows a 6.37% improvement. Other configurations, such as adding original books and newspapers or nonfiction books, also demonstrate positive gains of 5.51% and 3.22%, respectively. Conversely, the addition of fiction books is the only one to show a negative performance, with a decrease of 1.40%. Interestingly, when decomposed by skill, the addition of fiction books makes the model excel at generating more diverse texts.
@@ 6.3 Instruction-tuned Models
Lastly, as shown in Figure 5, when the core models are further fine-tuned on data to follow instructions, the gains across models are all consistent, showing that the core advantage lies in the pre-training data, while further training on instructions gives a consistent boost in performance. Moreover, instruction tuning also seems to reduce the gap between the base and extended configurations, suggesting that copyrighted material might become less relevant as supervised finetuning datasets increase in size in the post-training phases of LLMs. Interestingly, adding Norwegian instruction data on top of the base model seems enough to improve over the performance of Mistral 7B v0.1.
@@ 7 Discussion
Our findings underline the value of copyrighted materials in improving the performance of generative language models, particularly for specialized NLP tasks in Norwegian. The inclusion of curated, high-quality texts provides a substantial advantage in terms of language richness, coherence, and context-specific understanding. At the same time, these advantages are much less pronounced for the models warm-started with the weights pre-trained on other languages (mostly English). We see two possible reasons for this:
@@
    Copyrighted Norwegian data is indeed beneficial for LLMs, but the original models used for warm-starting were presumably already pre-trained on parts of this data. The exact composition of the training sets of models like Mistral is not known; it is quite probable that they may contain copyrighted content. Thus, continuous pre-training on (partially) the same content did not bring the expected benefits to the warm-started extended models.
@@ 7.2 Implications for Policy
The empirical evidence gathered in our research is crucial for informing copyright policy in the digital age. Policymakers can use these findings to establish frameworks that ensure creators are adequately compensated, balancing the needs of LLM innovation with the rights of authors and publishers. This is particularly relevant in light of ongoing lawsuits against major AI companies, underscoring the necessity of legal clarity.
@@ 8 Conclusion
Our study represents a pioneering effort to quantify the impact of copyrighted materials on LLMs for Norwegian. Our results indicate that high-quality copyrighted content significantly enhances model performance, especially for complex NLP tasks. However, these benefits bring forth ethical and legal challenges that must be addressed to ensure a sustainable and fair approach to LLM development. By providing empirical evidence, we aim to contribute to the ongoing discourse on the role of copyright in AI and inform future policies that support both innovation and the rights of content creators.
@@ 9 Future Work
Future work should focus on testing models at various scales and different pre-trained open weights to better understand how dataset composition affects performance. By experimenting with models of different sizes, we could identify any scaling thresholds where the impact of copyrighted material varies significantly.
Additionally, the observed effects of fiction on model performance highlight the need to 1) examine how different types of fiction—such as fantasy or historical fiction—impact tasks like Sentiment Analysis and Commonsense Reasoning, and 2) design new and adequate benchmarks for evaluating the contribution of fiction in Norwegian LLMs for tasks such as creative writing, plot understanding, or descriptive language use. This targeted investigation could clarify the role of fiction in model training and help refine data curation strategies.
@@ 10 Distribution
The base dataset and models were intended to be freely distributed, as all materials included were granted redistribution permissions under different agreements. After we communicated the results of our investigations to the different partners, some rightholders demanded a reinterpretation of the agreements in the light of the results and this new era of generative AI. This will most likely prohibit us from sharing publicly the exact models trained within the Mímir project, but we will publish a Norwegian language model trained on a version of the base dataset with this material excluded.

### v2

@@ 6 Results
The evaluation of the trained models demonstrated that incorporating publisher-controlled copyright-protected corpora provided a measurable boost in performance across a range of NLP tasks. To illustrate the overall performance differences, Figure 2 shows the total scores across all skills, averaged by task for each model. Non-aggregated scores for all tasks, prompts, and models are available at the Mímir repository. [^18] https://github.com/mimir-project/mimir-evaluation
@@ 6.1 Core Models
As shown in Table 4 and Figure 2, the performance analysis of core models across various tasks reveals distinct strengths for different configurations. The base (warm-started) configuration consistently excels in Sentiment Analysis, World Knowledge, and Norwegian Language. In contrast, the extended (warm-started) configuration leads in Fairness & Truthfulness, Reading Comprehension, Commonsense Reasoning, Translation, and Variation & Readability, indicating its robust performance for language-intensive tasks. The base configuration generally lags behind others, scoring the lowest across multiple tasks. Meanwhile, the extended configuration performs well, particularly in Summarization. Furthermore, it indicates that we could leverage the existing metadata available at the National Library to tailor subsets of the publisher-controlled copyrighted corpora and build models that excel at specific tasks. However, the difference between the base and extended warm-started models is very small. Further testing is required to assess whether this difference is still statistically significant [^19] Detailed scores available in Appendix F Table 8..
@@ 6.2 Domain-tuned Models
To further explore the specific effects of different types of training data, we analyzed the gains in performance by focusing on different sub-corpora, such as newspapers, books, and mixed datasets. Figure 3 provides an overview of the average percentage gains for models trained on various data configurations compared to the base model. It shows that the extended model exhibits the highest average gain at 6.73%, indicating substantial overall improvement. The addition of nonfiction books and newspapers follows with a 6.52% gain, and the addition of only newspapers shows a 6.37% improvement. Other configurations, such as adding original books and newspapers or nonfiction books, also demonstrate positive gains of 5.51% and 3.22%, respectively. Conversely, the addition of fiction books is the only one to show a negative performance, with a decrease of 1.40%. Interestingly, when decomposed by skill, the addition of fiction books makes the model excel at generating more diverse texts (see Figure 5 in Appendix E).
@@ 6.3 Instruction-tuned Models
Lastly, as shown in Figure 4, when the core models are further fine-tuned on data to follow instructions, the gains across models are all consistent, showing that the core advantage lies in the pre-training data, while further training on instructions gives a consistent boost in performance. Instruction tuning also seems to reduce the gap between the base and extended configurations, suggesting that publisher-controlled copyrighted corpora might become less relevant as supervised fine-tuning datasets increase in size in the post-training phases of LLMs. Interestingly, adding Norwegian instruction data on top of the extended model seems enough to improve over the performance of Mistral 7B v0.1.
@@ 7 Discussion
Our findings underline the value of copyrighted materials in improving the performance of generative language models, particularly for specialized NLP tasks in Norwegian. The inclusion of these curated publisher-controlled texts provide a substantial advantage in terms of language richness, coherence, and context-specific understanding. However, these advantages are significantly less evident in models that are warm-started using weights pre-trained on other languages, primarily English. We see two possible reasons for this:
@@
    Publisher-controlled copyrighted Norwegian data is indeed beneficial for LLMs, but the original models used for warm-starting were presumably already pre-trained on datasets that may share similarities with this data. Due to the lack of transparency regarding the exact composition of training datasets in models like Mistral, concerns about potential data contamination remain relevant. This overlap could explain why continuous pre-training on similar content did not yield the expected benefits for the warm-started extended models Li et al. (2024); Dong et al. (2024); Xu et al. (2024); Samuel et al. (2024).
@@ 7.2 Implications for Policy
The empirical evidence gathered in our research is crucial for informing copyright policy in the digital age. Policymakers can use these findings to establish frameworks that ensure creators are adequately compensated, balancing the needs of LLM innovation with the rights of authors and publishers. This is particularly relevant in light of ongoing lawsuits against major AI companies.
@@ 8 Conclusion
Our study represents a pioneering effort to quantify the impact of copyrighted materials on LLMs for Norwegian. Our results indicate that high-quality publisher-controlled copyrighted content significantly enhances model performance, especially for complex NLP tasks. However, these benefits bring forth ethical and legal challenges that must be addressed to ensure a sustainable and fair approach to LLM development. By providing empirical evidence, we aim to contribute to the ongoing discourse on the role of copyright in AI and inform future policies that support both innovation and the rights of content creators.
@@ 9 Future Work
Future work should focus on testing models at various scales and different pre-trained open weights to better understand how dataset composition affects performance. By experimenting with models of different sizes, we could identify any scaling thresholds where the impact of copyrighted material varies significantly. In retrospect, one notable flaw in the experimental design is the lack of fully traceable and transparent models, such as OLMo Groeneveld et al. (2024), which provide detailed documentation of their training data and processes. Without utilizing models with verifiable data provenance, it becomes challenging to accurately assess how specific dataset compositions, including copyrighted or genre-specific materials, influence model behavior and performance for warm-started models. Incorporating traceable models would improve the reproducibility and reliability of findings, ensuring that conclusions drawn about the impact of various text genres are well-founded.
Additionally, the observed effects of fiction on model performance highlight the need to 1) examine how different types of fiction –such as fantasy or historical fiction– impact tasks like Sentiment Analysis and Commonsense Reasoning, and 2) design new and adequate benchmarks for evaluating the contribution of fiction in Norwegian LLMs for tasks such as creative writing, plot understanding, or descriptive language use. This targeted investigation could clarify the role of fiction in model training and help refine data curation strategies.
@@ 10 Distribution
The base dataset and models were intended to be freely distributed, as all materials included were granted redistribution permissions under different agreements. After we communicated the results of our investigations to the different partners, some right-holders demanded a reinterpretation of the agreements (primarily the Language Technology Use, Språkteknologiformål), in the light of the results and this new era of generative AI. This prevented us from sharing publicly the exact models trained within the Mímir project, but instead we built a subset of base, which we are calling core, excluding the affected newspapers (around 1B words) and trained models both from scratch and from Mistral 7B v0.1. Their performance is on par with their base counterparts. We are also releasing these models under a permissive license [^20] https://huggingface.co/mimir-project/mimir-mistral-7b-core-scratch and https://huggingface.co/mimir-project/mimir-mistral-7b-core.
Acknowledgments
We extend our sincere gratitude to Hans Eide from Sigma2 for facilitating access to the LUMI supercomputer, enabling the computationally intensive tasks integral to this study. Additionally, we thank Google for providing compute resources via the Tensor Research Cloud program, which significantly supported our model training efforts.
This project would not have been possible without the trust and collaboration of the Ministry of Culture, which empowered the National Library of Norway to spearhead this endeavor, with the invaluable contributions of the Norwegian University of Science and Technology (NTNU) and the University of Oslo (UiO), whose expertise and insights were instrumental throughout the process. We are grateful for their vision and faith in the potential of this research.
We are also deeply appreciative of Olaus Bergstrøm and the entire legal team at the National Library of Norway for their guidance on the legal dimensions of this research. Their expertise was invaluable in navigating the complexities of copyright law and ensuring compliance with the unique considerations surrounding the materials used in this project.
A special thanks goes to the representatives of the Norwegian rights-holder organizations, who not only agreed to the use of their materials for this project but were steadfast in their support of the initiative. Their cooperation and encouragement have been vital in ensuring the project’s success and advancing research on the intersection of copyright and AI development.
Appendices
Appendix A Legal Cases
  • Bartz v. Anthropic PBC, No. 3:24-cv-05417 (N.D. Cal. Aug. 19, 2024)
  • The Ctr. for Investigative Reporting v. OpenAI, Inc., No. 1:24-cv-04872 (S.D.N.Y. Jun. 27, 2024)
  • UMG Recordings, Inc. v. Uncharted Labs, LLC, No. 1:24-cv-04777 (S.D.N.Y. Jun. 24, 2024)
  • UMG Recordings, Inc. v. Suo, Inc., No. 1:24-cv-11611 (D. Mass. Jun. 24, 2024)
  • J. L. v. Alphabet, Inc., No. 3:23-cv-03440, 2024 WL 3282528 (N.D. Cal. June 6, 2024)
  • In re OpenAI ChatGPT Litigation, No. 3:23-cv-03223, 2024 WL 2044625 (N.D. Cal. May 7, 2024)
  • Makkai v. Databricks, Inc., No. 4:24-cv-02653 (N.D. Cal. May 2, 2024)
  • Dubus v. NVIDIA Corp., No. 3:24-cv-02655 (N.D. Cal. May 2, 2024)
  • Daily News LP v. Microsoft Corp., No. 1:24-cv-03285 (S.D.N.Y. Apr. 30, 2024)
  • Zhang v. Google LLC, No. 3:24-cv-02531 (N.D. Cal. Apr. 26, 2024)
  • Nazemian v. NVIDIA Corp., No. 4:24-cv-01454 (N.D. Cal. Mar. 8, 2024)
  • The Intercept Media, Inc. v. OpenAI, Inc., No. 1:24-cv-01515 (S.D.N.Y. Feb. 28, 2024)
  • Raw Story Media, Inc. v. OpenAI Inc., No. 1:24-cv-01514 (S.D.N.Y. Feb. 28, 2024)
  • Tremblay v. OpenAI, Inc., No. 3:23-cv-03233, 2024 WL 557720 (N.D. Cal. Feb. 12, 2024)
  • Universal Music Group v. Anthropic (February 2024)
  • The New York Times v. OpenAI & Microsoft (December 2023)
  • Kadrey v. Meta Platforms, Inc., No. 3:23-cv-03417, 2023 WL 10673221 (N.D. Cal. Dec. 1, 2023)
  • Alter v. OpenAI Inc., No. 1:23-cv-10211 (S.D.N.Y. Nov. 21, 2023)
  • Andersen v. Stability AI Ltd., No. 3:23-cv-00201, 2023 WL 7132064 (N.D. Cal. Oct. 30, 2023)
  • Concord Music Group, Inc. v. Anthropic PBC, No. 3:23-cv-01092 (M.D. Tenn. Oct. 18, 2023)
  • Huckabee v. Meta Platforms, Inc., No. 3:23-cv-06663 (N.D. Cal. Oct. 17, 2023)
  • Authors Guild v. OpenAI Inc., No. 1:23-cv-08292 (S.D.N.Y. Sept. 18, 2023)
Appendix B Sampling
We built three custom perplexity models for specific Norwegian domains that proved too divergent from Wikipedia: books, newspapers, and government documents. These perplexity models were used to score each document in the datasets. Based on their perplexity scores, the documents were further divided into three segments corresponding to their quartile distribution. Documents with scores below the first quartile were classified as “good”, those between $Q_{1}$ and $Q_{3}$ as “medium”, and those above $Q_{3}$ were considered “bad”. The documents in each segment were randomized. While the intention was to train all models on progressively better data, starting from “bad” segment, then “medium” and finally the “good” segment, we never got around to test whether this approach would result in better performing models.
Moreover, from the clean and deduplicated sources, we sub-sampled each non-Norwegian language at an specific sampling ratio until achieving the proportion of documents shown in Figure 5. Pseudo-code for the algorithm used to subsample is shown in Algorithm 1212121https://huggingface.co/mimir-project/mimir-perplexity. We also discovered that a good amount of documents were misclassified by the fastText language identifier Joulin et al. (2016).
Language Sampling ratio Final ratio
Bokmål 100.00% 35.74%
Danish 43.00% 8.01%
English 81.00% 4.53%
Icelandic 100.00% 1.31%
Nynorsk 100.00% 2.02%
Swedish 15.40% 4.46%
Code 62.00% 4.53%
Table 5: Percentage of documents kept from the clean and deduplicated sources and the final proportion of documents in each language present in the final dataset. Code was considered its own language when sampling.
Algorithm 1 Sub-sampling Dataset Based on Perplexity Distribution
1:Input: Dataset $D$ with perplexity distribution, target sampling ratio $R$
2:Output: Sub-sampled dataset $D^{\prime}$
3:procedure Subsample($D,R$)
4:     Compute the quartile values $q_{1}$ and $q_{3}$ from the perplexity distribution of $D$
5:     Define an initial Gaussian PDF with mean $\mu=(q_{1}+q_{3})/2$ and standard deviation $\sigma$ such that $q_{1}$ and $q_{3}$ align with the corresponding positions in the Gaussian curve
6:     Compute the histogram $H$ of perplexity values from $D$
7:     Combine $H$ with the Gaussian weights to estimate the initial sampling ratio $R_{0}$
8:     Compute the normalization factor $N$ such that $R_{0}\times N=R$
9:     while Error in central quartile probabilities exceeds tolerance do
10:         Adjust the parameters $\mu$ and $\sigma$ of the Gaussian curve to minimize the error in the desired probabilities within the central quartiles $[q_{1},q_{3}]$
11:         Update the normalization factor $N$ to match the target ratio $R$
12:     end while
13:     for each sample $s$ in $D$ do
14:         Compute the perplexity $p_{s}$ of sample $s$
15:         Estimate the probability $P(s)$ of retaining sample $s$ based on the normalized Gaussian PDF
16:         if $P(s)\geq$ random threshold then
17:              Retain $s$ in the sub-sampled dataset $D^{\prime}$
18:         end if
19:     end for
20:end procedure
21:return $D^{\prime}$
Appendix C Sources
Source Raw Clean extended base
Books 3.7B 2.5B 1.9B 1.9B
CulturaX 52.7B 52.1B 21.8B 16.9B
Digimanus 9.6M 4.6M 3.4M 3.3M
Evaluerings- rapport 76.7M 68.6M 61.2M 61.5M
HPTL v1.2 35.5B 34.1B 14.9B 11.3B
LovData 57.1M 57.1M 53.7M 54.8M
Målfrid 7.5B 1.9B 1.7B 1.7B
Newspapers 4.6B 3.6B 3.2B 3.3B
Parlamint 170.3M 84.4M 83.4M 83.3M
PG19 2.0B 1.9B 1.4B 428.6M
StarCoder 19.7B 9.8B 7.1B 3.4B
Wikipedia 4.0B 3.9B 2.8B 996.2M
Books (restricted) 21.7B 20.0B 18.1B 0
Newspapers (restricted) 14.3B 9.8B 9.1B 0
Total 166.1B 139.8B 82.1B 40.1B
Table 6: Number of words (comma separated) per source at the start of the data pipeline (raw count), after cleaning, and in the extended and base datasets.
Appendix D Hyperparameters
Hyperparameter Core Models Domain-Tuned Models Instruction-tuned Models
Model size 7B 7B 7B
Hidden layers 32 32 32
Attention heads 32 32 32
Hidden size 4096 4096 4096
Intermediate size 14336 14336 14336
Max position embeddings 2048 2048 2048
Key-value heads 8 8 8
Sliding window 4096 4096 4096
Precision bfloat16 bfloat16 bfloat16
Optimizer AdamW AdamW AdamW
Optimizer parameters $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$ $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$ $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$
Global batch size 4M ($2048\times 2048$) tokens 4M ($2048\times 2048$) tokens 512 seqs
Initial/final learning rate $3.0\times 10^{-4}$ / $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ / $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ / $3.0\times 10^{-7}$
Vocabulary size 32768 32768 32768
Training steps 64k 10k 4 epochs
Dropout 0 0 0
Warm-up steps 2000 200 20
Weight decay 0.1 0.1 0.1
Checkpoints Every 1000 steps Every 1000 steps Every 1 epoch
Shuffle Shuffle after each epoch Shuffle after each epoch Shuffle after each epoch
Table 7: Hyperparameters for the Mímir model set.
Appendix E Percentage Gains
Figure 5 illustrates the percentage gains of each domain-tuned model with respect to the performance of the base model, per higher level skill. Training on different materials shows distinct trade-offs: newspaper data excels at Translation (27.20% gain) and Norwegian Language (51.92%), while fiction books improve Variation & Readability (7.83%). Combining books and newspapers often yields balanced improvements, though most configurations struggle with Reading Comprehension and Translation. The extended configuration, which supplements books and newspapers with Internet data, shows strong all-around performance, particularly in Summarization (26.37%) and World Knowledge (5.60%).
basebase + translated booksbase + fiction booksbase + booksbase + nonfiction booksbase + newspapersbase + nonfiction books + newspapersbase + original books + newspapersbase + books + newspapersbase + original booksextended10.86%8.52%13.52%11.77%13.59%10.11%9.72%9.58%6.64%3.85%Sentiment Analysisbase-4.67%3.59%-0.84%-10.23%-10.33%-9.90%1.71%0.74%5.60%3.06%Fairness & Truthfulnessbase3.52%-13.59%-7.75%-6.89%-3.63%-7.22%-16.30%-12.24%-11.37%-9.20%Reading Comprehensionbasebase + translated booksbase + fiction booksbase + booksbase + nonfiction booksbase + newspapersbase + nonfiction books + newspapersbase + original books + newspapersbase + books + newspapersbase + original booksextended-3.32%-0.31%-0.11%5.60%1.13%1.05%1.48%1.82%1.08%1.08%World Knowledgebase-5.18%-0.65%3.55%3.08%4.81%4.04%7.09%2.04%3.86%8.72%Commonsense Reasoningbase24.43%38.55%36.13%40.57%42.37%51.92%37.10%35.03%7.56%13.98%Norwegian Languagebasebase + translated booksbase + fiction booksbase + booksbase + nonfiction booksbase + newspapersbase + nonfiction books + newspapersbase + original books + newspapersbase + books + newspapersbase + original booksextended-11.43%-1.80%-5.47%-5.13%-9.81%-9.01%-8.13%-12.43%-15.21%26.37%Summarizationbase-3.25%-6.47%-3.72%-2.18%-2.52%-6.42%-6.90%-10.24%-11.01%27.20%Translationbase1.99%1.96%2.68%2.15%2.41%1.98%0.80%2.65%7.83%4.16%Variation & Readability
Figure 5: Percentage gains over the performance of the base model per skill.
Appendix F Evaluation Scores
Model SA FT RC WK CR NL S T VR Score
Core Models
base 69.54 53.51 38.04 41.10 39.85 38.28 32.45 41.55 59.66 413.98
extended 77.09 51.01 39.38 43.40 41.26 47.64 41.00 40.20 60.85 441.83
base (warm) 88.17 51.28 52.48 53.27 48.02 49.51 35.88 41.14 60.52 480.28
extended (warm) 86.57 54.64 52.79 51.51 49.25 48.48 36.14 42.48 61.12 482.98
Domain Tuned Models
base + books 76.20 55.15 33.38 41.54 39.59 51.69 29.81 38.68 61.24 427.30
base + newspapers 76.57 48.04 35.29 41.55 37.79 58.16 29.26 52.85 60.85 440.35
base + books + newspapers 78.94 53.06 35.09 41.53 41.77 52.11 31.86 40.00 61.27 435.64
base + fiction books 74.16 47.99 33.71 39.74 41.39 41.18 28.41 37.29 64.33 408.20
base + nonfiction books 76.30 56.51 31.84 41.06 40.66 52.48 29.52 38.88 60.14 427.40
base + nonfiction books + newspapers 78.99 53.91 36.66 41.85 42.68 54.50 30.78 40.50 61.10 440.97
base + original books 75.46 55.43 32.87 41.56 41.08 53.04 28.74 38.86 60.83 427.87
base + original books + newspapers 77.72 54.43 35.42 41.71 41.46 53.81 30.67 40.64 60.95 436.81
base + translated books 72.22 48.21 34.54 40.97 43.33 43.63 27.51 36.97 62.15 409.54
Instruction Fine Tuned Models
base (warm) instruct 87.83 53.70 50.33 54.98 49.42 59.53 38.36 49.75 59.46 503.36
extended (warm) instruct 89.81 57.80 51.69 53.09 49.76 55.91 37.75 47.72 60.35 503.89
base instruct 69.45 50.59 36.27 41.18 42.06 53.53 35.14 58.83 58.35 445.39
extended instruct 78.90 46.10 38.68 44.32 43.57 56.46 36.40 54.64 59.29 458.36
Mistral 7B v0.1 88.41 64.93 56.68 58.86 36.01 31.49 10.09 41.55 59.99 448.01
Table 8: Detailed scores across all skills for each model configuration. Abbreviations: SA = Sentiment Analysis, FT = Fairness & Truthfulness, RC = Reading Comprehension, WK = World Knowledge, CR = Commonsense Reasoning, NL = Norwegian Language, S = Summarization, T = Translation, VR = Variation & Readability. Best overall scores per skill in bold. Best score per skill and model group underlined. Mistral 7B v0.1 also added for reference.